OpenStudy (anonymous):

what is 9/45 in lowest terms

OpenStudy (misty1212):

HI again!!

OpenStudy (er.mohd.amir):

9=3*3 and 45=3*3*5 cancel same number.

OpenStudy (misty1212):

there is a common factor of 9 top and bottom, divide top and bottom by 9

OpenStudy (anonymous):

thanks